# Pattern Analogies: Learning to Perform Programmatic Image Edits by Analogy

### v1

@@ 1 Introduction
Editing pattern images is not straightforward, as patterns are inherently structured, defined by rules that govern their layout and composition: tiling patterns adhere to principles of alignment and repetition (see Figure 1: top left), while retro-style designs rely on spatial divisions and fills (see Figure 1: bottom left). The edits that designers desire often aim to adjust these underlying organizational rules rather than make superficial, pixel-level changes. We refer to such edits as programmatic edits, requiring manipulation of the underlying program that defines a pattern’s structure.
@@
To make our approach possible, we introduce SplitWeave: a domain-specific language (DSL) for crafting visual patterns. SplitWeave serves two purposes in our method. First, it enables parametric definition of input pairs $(A,A^{\prime})$, allowing users to guide transformations in $(B,B^{\prime})$ as if the underlying program for $B$ were accessible. In Figure 1 (right), modifying the SplitWeave program for $A^{\prime}$ produces corresponding changes in $B^{\prime}$. Second, SplitWeave supports the creation of large-scale synthetic training data. We develop program samplers that generate high-quality patterns in two common styles: tiling-based designs with repeating elements and color field patterns characterized by splitting the canvas into intricate colored regions. Training a model for analogical editing requires a dataset of quartets $(A,A^{\prime},B,B^{\prime})$. By applying identical programmatic edits to the SplitWeave programs for both $A$ and $B$ to produce $A^{\prime}$ and $B^{\prime}$, we ensure that the transformation from $A$ to $A^{\prime}$ mirrors that from $B$ to $B^{\prime}$. This approach allows us to generate a diverse dataset of analogical quartets. Models trained on this dataset can generalize effectively to real-world patterns within these styles and can extend to related styles.
@@
To evaluate our method, we curated a test set of $50$ patterns from Adobe Stock spanning $7$ distinct styles. A perceptual study on this dataset shows that participants prefer edits by TriFuser over recent training-free and training-based methods. Although our training data covers only two of these styles, our model demonstrates effective generalization to the other, out-of-distribution styles. On a synthetic validation set with ground-truth analogical edits, our model produces outputs more similar to the ground truth than other methods. Finally, we showcase two applications of our approach: mixing attributes of different patterns and transferring pattern animations.
@@ 3.3 Learning an Analogical Editor
As we demonstrate in Section 4.4, conditioning on $C^{\Omega}$ instead of $C$ significantly enhances the quality of patterns generated by TriFuser. Our adapted architecture, shown in Figure 5, integrates the modifications described above to effectively address the described drawbacks. To enhance generalizability to real-world patterns, we initialize TriFuser with an existing pretrained IM model [65], and fine-tune only the denoising UNet and the projection layers in our feature extractor.
@@ 4.2 Analogical Editing Baselines
TriFuser vs. Analogist $87.74\%$
TriFuser vs. LatentMod $80.78\%$
TriFuser vs. Inpainter $72.21\%$
@@ 4.3 Editing Real-World Patterns
We performed a two-alternative forced-choice perceptual study comparing TriFuser with baseline methods on all 50 entries in the test set. Each method generates $k=9$ outputs for each input tuple, and we select the best one based on visual inspection. Participants were shown edited patterns generated by two different methods along with the input patterns $(A,A^{\prime},B)$ and instructed to select the edit that best preserved the analogical relationship and exhibited higher image quality. We recruited 32 participants for the study, resulting in a total of 1550 total judgments.
Table 1 presents the results, showing that TriFuser was preferred over both Analogist and LatentMod. Due to the domain gap between the training data of the underlying model [52] and pattern images, Analogist fails to interpret and edit pattern images. Meanwhile, LatentMod fails to perform reasonable edits as the embedding space lacks the low-level details necessary for programmatic edits While these baselines perform adequately on stylistic edits, they are unsuitable for programmatic editing. When compared to Inpainter, TriFuser was favored in 72.2% of comparisons. Both methods benefit from training on analogical quartets, yet Inpainter sacrifices pattern quality as it generates the edited pattern in only a quarter of the full canvas resolution.

### v2

@@ 1 Introduction
Editing pattern images is not straightforward, as patterns are inherently structured, defined by rules that govern their layout and composition: tiling patterns adhere to principles of alignment and repetition (see Figure 1: top left), while retro-style designs rely on spatial divisions and fills (see Figure 1: bottom center). The edits that designers desire often aim to adjust these underlying organizational rules rather than make superficial, pixel-level changes. We refer to such edits as programmatic edits, requiring manipulation of the underlying program that defines a pattern’s structure.
@@
To make our approach possible, we introduce SplitWeave: a domain-specific language (DSL) for crafting visual patterns. SplitWeave serves two purposes in our method. First, it enables parametric definition of input pairs $(A,A^{\prime})$, allowing users to guide transformations in $(B,B^{\prime})$ as if the underlying program for $B$ were accessible. In Figure 1 (right), modifying the SplitWeave program for $A^{\prime}$ produces corresponding changes in $B^{\prime}$. Second, SplitWeave supports the creation of large-scale synthetic training data. We develop program samplers that generate high-quality patterns in two common styles: tiling-based designs with repeating elements and color field patterns characterized by splitting the canvas into intricate colored regions. Training a model for analogical editing requires a dataset of quartets $(A,A^{\prime},B,B^{\prime})$. By applying identical programmatic edits to the SplitWeave programs for both $A$ and $B$ to produce $A^{\prime}$ and $B^{\prime}$, we ensure that the transformation from $A$ to $A^{\prime}$ mirrors that from $B$ to $B^{\prime}$. This approach allows us to generate a diverse dataset of analogical quartets. Models trained on this dataset generalize effectively to real-world patterns within these styles and even extend to related styles.
@@
To evaluate our method, we curated a test set of $50$ patterns from Adobe Stock spanning $7$ distinct styles. A perceptual study on this dataset reveals that participants consistently prefer edits by TriFuser over those produced by recent training-free and training-based analogical editing methods. Despite training only on two of these styles, our model successfully generalizes to the remaining five, which were unseen during training. On a synthetic validation set with ground-truth analogical edits, our model achieves higher structural and perceptual similarity to the ground truth than prior methods Finally, we showcase two compelling applications of our approach: mixing attributes of different patterns and transferring pattern animations.
@@ 3.3 Learning an Analogical Editor
As we demonstrate in Section 4.4, conditioning on $C^{\Omega}$ instead of $C$ significantly enhances the quality of patterns generated by TriFuser. Our adapted architecture, shown in Figure 5, integrates the modifications described above to effectively address the described drawbacks. Note that this architecture is designed for general analogical editing, not just pattern editing. To enhance generalizability to real-world patterns, we initialize TriFuser with an existing pretrained IM model [65], and fine-tune only the denoising UNet and the projection layers in our feature extractor.
@@ 4.2 Analogical Editing Baselines
TriFuser vs. Analogist $89.24\%$
TriFuser vs. LatentMod $80.59\%$
TriFuser vs. Inpainter $72.05\%$
@@ 4.3 Editing Real-World Patterns
We performed a two-alternative forced-choice perceptual study comparing TriFuser with baseline methods on all 50 entries in the test set. Each method generates $k=9$ outputs for each input tuple, and we select the best one based on visual inspection. Participants were shown edited patterns generated by two different methods along with the input patterns $(A,A^{\prime},B)$ and instructed to select the edit that best preserved the analogical relationship and exhibited higher image quality. We recruited 42 participants for the study, resulting in a total of 1720 total judgments.
Table 1 presents the results, showing that TriFuser was preferred over both Analogist and LatentMod. Due to the domain gap between the training data of the underlying model [52] and pattern images, Analogist fails to interpret and edit pattern images. Meanwhile, LatentMod fails to perform reasonable edits as the embedding space lacks the low-level details necessary for programmatic edits While these baselines perform adequately on stylistic edits, they are unsuitable for programmatic editing. When compared to Inpainter, TriFuser was favored in 72.05% of comparisons. Both methods benefit from training on analogical quartets, yet Inpainter sacrifices pattern quality as it generates the edited pattern in only a quarter of the full canvas resolution.